Business Weekly conducted a survey of graduates from 30 top MBA programs. On the basis of the survey, assume the mean annual salary for graduates 10 years after graduation is 162000 dollars. Assume the population standard deviation is 38000 dollars. Suppose you take a simple random sample of 96 graduates. Find the probability that a single randomly selected salary that doesn't exceed 166000 dollars.

Answer = ______

Find the probability that a sample of size n = 96 is randomly selected with a mean that that doesn't exceed 166000 dollars.

Answer =______

Enter your answers as numbers accurate to 4 decimal places.
Answer #1

Solution :

Given that ,

a) P(x < 166000)

= P[(x - \mu ) / \sigma < (166000 - 162000) / 38000]

= P(z < 0.105)

Using z table,

= 0.5418

b) \mu\bar x = 162000

\sigma\bar x = \sigma / \sqrt n = 38000 / \sqrt 96

P(\bar x < 166000) = P((\bar x - \mu \bar x ) / \sigma \bar x < (166000 - 162000) / 38000 / \sqrt 96 )

= P(z < 1.031)

Using z table

= 0.8487